OpenStudy (micheal1234):

what is the solution of m over 4-3 =7

OpenStudy (anonymous):

are you trying to find m? so m/1 =7?

OpenStudy (anonymous):

can you attach link plz

OpenStudy (anonymous):

It seems like it's \[\frac{ m }{ 4 }-3=7\] Add the 3 over m/4=10 Then solve for m with simple multiplication.

OpenStudy (micheal1234):

yes

OpenStudy (anonymous):

40

OpenStudy (dayakar):

\[\frac{ m }{ 4 }-1=7\] \[\frac{ m }{ 4 }-1+1=7+1\] adding bothsides 1 \[\frac{ m }{ 4 }=8\] 4\[4\times \frac{ m }{ 4 }=4\times8\] multiplying bothsides with 4 therefore m= 32

OpenStudy (anonymous):

Where do you see that equation? @dayakar

OpenStudy (johnn.):

One does not equal seven.

OpenStudy (anonymous):

to balance the variables out he's not saying they equal and its showing that -1=-1